# Philippe Brax    Pierre Vanhove CEA, DSM, Institut de Physique Théorique, IPhT, CNRS, MPPU, URA2306, Saclay, F-91191 Gif-sur-Yvette, France

### v1

@@ I Introduction
where we have introduced $\Omega_{\Lambda 0}\sim 0.7$ as the fraction of dark energy in the Universe and $\rho_{c}=3M_{\textrm{Pl}}^{2}H_{0}^{2}$ is the critical density of the Universe.
@@
can be estimated using the Planck data to be a large contribution
@@
which exemplifies the nature of the cosmological constant problem even at low energy.
@@
This reasoning makes use of the presence at low redshift of hot high-density structures within the cooler and lower-density cosmological background.
@@ II The $\mathcal{R}^{2}$ gravity
The contributions $\mathcal{L}_{\rm corrections}$ arise from such extensions of Einstein gravity induced either by high-energy quantum corrections from a high-energy completion like string theory or new interactions from extra massless fields which are still undetected.
@@
it has been shown in Fradkin:1981iu ; Salvio:2014soa that the coefficient $c_{2}(\mu)$ is always asymptotically free, since $dc_{2}(\mu)/d\log\mu^{2}>0$, whereas $c_{0}(\mu)$ is asymptotically safe, $dc_{0}(\mu)/d\log\mu^{2}<0$ for the non tachyonic case $c_{0}(\mu)>0$, which is the case of interest in this paper. Therefore, at low-energy $c_{2}(\mu)$ tends to zero whereas $c_{0}(\mu)$ grows, leading to the hierarchy $c_{0}(\mu)\gg c_{2}(\mu)$ at very low energy. Hence the quadratic Ricci scalar term is enhanced as compared with other quadratic and higher order contributions.
@@ III $\mathcal{R}^{2}$ running from inflation to Dark Energy
The initial values of the renormalisation group are taken at the end of inflation corresponding to the reheat temperature $T_{\rm reh}$, i.e. larger than any physical masses of the particles in the spectrum of the theory.
@@ IV Observational predictions
We remark that this interval is compatible with the value $m_{\rm scalaron}\simeq 4.4\times 10^{-3}$ eV for which the $\mathcal{R}^{2}$ induced scalaron could be at the origin of the observed dark matter abundance Cembranos:2008gj ; Cembranos:2015svp ; Shtanov:2021uif ; Shtanov:2022xew . In this scenario, at low energy compared to the inflation scale, the vacuum expectation value of the scalaron is displaced from the origin by an amount depending on the electroweak scale $v\simeq 250$ GeV because of its coupling to the Higgs field. For more generic initial conditions after inflation taking into account the quantum fluctuations of the scalaron during inflation, the whole interval up to $m_{\rm scalaron}\simeq 0.1$ eV could accommodate dark matter. As the electroweak transition begins, the scalaron starts oscillating with a decreasing amplitude eventually converging to the origin. This misalignment mechanism is similar to what happens for axions and leads an abundance of dark matter which fits the observed value for $m_{\rm scalaron}\simeq 4.4\times 10^{-3}$ eV Shtanov:2022xew . Combining both scenarios, this would lead to a possible signal in gravitational experiments below a distance $d\lesssim 45~{}\mu{\rm m}$, which is well within the allowed interval (16). This strongly supports the exciting possibility of testing the existence of the new gravitational interaction ${\cal R}^{2}$, which could play a role in both the late acceleration of the universe and dark matter. All in all, we believe that this a is a consequence of the universality of the effective long distance physics of gravity compared to the scales of quantum gravity.

### v2

# Philippe Brax and Pierre Vanhove CEA, DSM, Institut de Physique Théorique, IPhT, CNRS, MPPU, URA2306, Saclay, F-91191 Gif-sur-Yvette, France

@@ I Introduction
where we have introduced $\Omega_{\Lambda 0}\sim 0.7$ as the fraction of dark energy in the Universe and $\rho_{c}=3M_{\textrm{Pl}}^{2}H_{0}^{2}$ is the critical density of the Universe. Here $H_{0}$ is the Hubble rate today and $M_{\textrm{Pl}}=\sqrt{\hbar c/(8\pi G_{N})}$ is the reduced Planck mass.
@@
can be estimated using the Planck data Planck:2018vyg to be a large contribution
@@
which exemplifies the nature of the cosmological constant problem even at low energy111 The Planck data constrains the sum of the neutrino masses $\sum_{\nu}m_{\nu}\leq 0.12{\rm eV}$ at 95 % confidence level..
@@
This is the Weinberg bound which guarantees that the existence of galaxies is not prevented by the cosmological constant. Larger values would prevent the existence of galaxies and therefore clusters. The factor of 200 is obtained using the spherical collapse approximation for a separate Universe and reads more precisely Weinberg:1987dv
$|\Lambda^{4}(m_{e})|\lesssim\frac{500}{729}\,\delta_{R}\rho_{R}.$ (7)
where $\rho_{R}$ is the matter density at recombination and $\delta_{R}$ the typical overdensity at the same epoch. This reasoning makes use of the presence at low redshift of hot high-density structures within the cooler and lower-density cosmological background.
@@ II The $\mathcal{R}^{2}$ gravity
The contributions $\mathcal{L}_{\rm corrections}$ arise from such extensions of Einstein gravity induced either by high-energy quantum corrections from a high-energy completion like string theory or new interactions from extra massless fields which are still undetected. For instance in a string context and for curvatures below the string scale, only the curvature corrections up to cubic order are relevant. Higher order corrections in the curvature are suppressed by powers involving the ratio with the string scale Burgess:2016owb ; Brax:2017bcp . Similarly the cubic term in the curvature is suppressed compared to the quadratic terms as long as the curvature is less than the compactification scale. In summary and as long as one considers scales below the string and compactification scales, models described by quadratic curvature corrections are sufficient. Of course this applies to the description of inflation as long as the inflation scale is low enough compared to the string and compactification scales.
@@
it has been shown in Julve:1978xn ; Fradkin:1981iu ; Avramidi:1985ki ; Salvio:2014soa that the coefficient $c_{2}(\mu)$ is always asymptotically free, since $dc_{2}(\mu)/d\log\mu^{2}>0$, whereas $c_{0}(\mu)$ is asymptotically safe, $dc_{0}(\mu)/d\log\mu^{2}<0$ for the non tachyonic case $c_{0}(\mu)>0$, which is the case of interest in this paper. Therefore, at low-energy $c_{2}(\mu)$ tends to zero whereas $c_{0}(\mu)$ grows, leading to the hierarchy $c_{0}(\mu)\gg c_{2}(\mu)$ at very low energy. Hence the quadratic Ricci scalar term is enhanced as compared with other quadratic and higher order contributions.
@@ III $\mathcal{R}^{2}$ running from inflation to Dark Energy
The initial values of the renormalisation group are taken at the end of inflation corresponding to the reheat temperature $T_{\rm reh}$, i.e. larger than any physical masses of the particles in the spectrum of the theory. The previous renormalisation group equations allow one to calculate the evolution of $c_{0}$ and $\Lambda$ from inflation to the infrared. The only new ingredient in this integration is the presence of phase transitions taking place at different values of $\mu$, for instance the QCD phase transition, where the cosmological constant jumps due to the change of vacuum energy during the transitions Bellazzini:2015wva .
@@ IV Observational predictions
We remark that this interval is compatible with the value $m_{\rm scalaron}\simeq 4.4\times 10^{-3}$ eV for which the $\mathcal{R}^{2}$ induced scalaron could be at the origin of the observed dark matter abundance Cembranos:2008gj ; Cembranos:2015svp ; Shtanov:2021uif ; Shtanov:2022xew . In this scenario, at low energy compared to the inflation scale, the vacuum expectation value of the scalaron is displaced from the origin by an amount depending on the electroweak scale $v\simeq 250$ GeV because of its coupling to the Higgs field. For more generic initial conditions after inflation taking into account the quantum fluctuations of the scalaron during inflation, the whole interval up to $m_{\rm scalaron}\simeq 0.1$ eV could accommodate dark matter. As the electroweak transition begins, the scalaron starts oscillating with a decreasing amplitude eventually converging to the origin. This misalignment mechanism is similar to what happens for axions and leads an abundance of dark matter which fits the observed value for $m_{\rm scalaron}\simeq 4.4\times 10^{-3}$ eV Shtanov:2022xew . Combining both scenarios, this would lead to a possible signal in gravitational experiments below a distance $d\lesssim 45~{}\mu{\rm m}$, which is well within the allowed interval (17) 222 The link with scalar dark matter is guaranteed as long as the amplitude of oscillation around the minimum of the scalar potential is small and higher order terms act as corrections to a simple quadratic oscillator. When the quadratic terms are dominant the scalar dark matter is of the fuzzy type Hui:2016ltb whilst corrections to the quadratic Lagrangian can also be interesting astrophysically Brax:2019fzb . We leave a detailled study of the phenomenology for future work. . This strongly supports the exciting possibility of testing the existence of the new gravitational interaction ${\cal R}^{2}$, which could play a role in both the late acceleration of the universe and dark matter. All in all, we believe that this a is a consequence of the universality of the effective long distance physics of gravity compared to the scales of quantum gravity.